OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. x - 4y = 1 2x - 9y = 3 (-1, -3) (-3, 1) (3, 1) (-3, -1)

OpenStudy (anonymous):

In the first equation, we solve for x: x = 1+4y. As the name suggests, we plug this x-expression into the second equation: 2(1+4y) -9y = 2 + 8y - 9y = 2 -y = 3 (by hypothesis). Thus, we conclude y = -1. Then, we know x = 1+4y, thus, x = -3. We have (-3,-1)